Circular No. 7370

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 D1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports observations of an evident Kreutz sungrazing comet with a tail discovered by D. Lewis in SOHO/LASCO C3 data. Astrometry and parabolic orbital elements are given on MPEC 2000-D38. Biesecker provides apparent magnitudes brightening from V = 7.4 ± 0.2 on Feb. 28.971 to 5.8 ± 0.1 on Feb. 29.404 UT.

2000 UT  α_{2000}  δ_{2000}
Feb. 28.571  23^{h}15.^{m}0  −09^{o}57′

COMET P/1999 XB_{69} (LINEAR) An apparently asteroidal object discovered by LINEAR (MPS 8018)

with a cometlike orbit has been observed by C. Hergenrother, Lunar and Planetary Laboratory, on Feb. 27 with the Catalina 1.54-m reflector to show a 5′′ coma and a 10′′ tail in p.a. 80^{o}. Additional astrometry and orbital elements (q = 2000 Feb. 17.04 TT, q = 1.64 AU, i = 11^{o}.34, P = 9.4 yr) appear on MPEC 2000-D39.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 7.29234  5^{h}06^{m}13.^{s}11  +22^{o}55′00′′.3  17.9

COMET P/1999 XN_{120} (CATALINA) An apparently asteroidal object discovered by the Catalina Sky Survey

(MPS 8071) with a cometlike orbit was also observed by Hergenrother on Feb. 27 with the 1.54-m reflector to show a 12′′ coma but no tail. Additional astrometry and orbital elements (q = 2000 May 1.41 TT, q = 3.29 AU, i = 5^{o}.03, P = 8.5 yr) appear on MPEC 2000-D40.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 5.19023  3^{h}28^{m}46.^{s}75  +24^{o}01′26′′.8  16.9

CORRIGENDA On IAUC 6956, ‘Supernova 1998cr in PGC 57979’, line 5, for 0′′.5 west

read 6′′.0 west

On IAUC 6957, ‘Supernova 1998cs in UGC 10432’, line 5, for 0′′.2 east read 2′′.2 east

2000 February 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green